OpenStudy (anonymous):

Will someone please explain this to me? A motorboat takes 5 hours to travel 100km going upstream. The return trip takes 2 hours going downstream. What is the rate of the boat in still water and what is the rate of the current?

OpenStudy (inowalst):

@Compassionate

OpenStudy (ganpat):

Let the speed Upstream be a kmph and in downstream be b kmph, a = 100/5 and b = 100/2 Speed in Still water = (a + b)/2 Rate of stream = (a - b)/2

OpenStudy (anonymous):

So speed in water would be 35 and rate of stream would be 15?

OpenStudy (ganpat):

according to the formula.. Yes !

OpenStudy (anonymous):

Thank you so much!

OpenStudy (ganpat):

Sure ! :)